Aga was 9 yrs. old. When his sister was born. Now, he is 6 yrs less than twice his sister's age. What are their present ages?
Answer by stanbon(75887)

Aga was 9 yrs. old. When his sister was born. Now, he is 6 yrs less than twice his sister's age. What are their present ages?
----
Equations:
a-s = 9
---
a = 2s-6
----
Substitute for "a" and solve for "s":
9+s = 2s-6
s = 15 yrs: (sister's age now)
----
Solve for a using a = s+9
a = 15+9 = 24yrs (Aga's age now)
===============
Cheers,
Stan H.
===============